The police department is concerning about the number of parking spaces in the campus. The calculations show that the risk of shortage in parking spaces appears if proportion of students driving their own cars increase to 35%. They conducted a student survey and found that 93 out of 368 students drive their own car. Find 95% confidence interval for proportion of all Texas State students who are driving their own car. Is there statistical evidence of the risk of parking shortage?

a. What is the population parameter to estimate?

b. What is the sample statistic provided?

c. Find a 95% confidence interval for population parameter.

d. Interpret your interval in the context.

e. Is there statistical evidence of the risk of parking shortage?
Answer #1

Solution-a:

population parameter to estimate=p=0.35

b. What is the sample statistic provided?

sample statistic,p^=x/n=93/368= 0.2527174

c. Find a 95% confidence interval for population parameter.

z crit for 95%=1.96

95% confidence interval for p is

p^-z*sqrt(p^*(1-p^)/n,p^+z*sqrt(p^*(1-p^)/n

0.2527174-1.96*sqrt(0.2527174*(1-0.2527174)/368),0.2527174+1.96*sqrt(0.2527174*(1-0.2527174)/368)

0.2083165, 0.2971183

d. Interpret your interval in the context.

we are 95% confident that the true population proportion of students driving their own cars lies oin between

0.2083165, 0.2971183